# ChIRAAG: ChatGPT Informed Rapid and Automated Assertion Generation

### v2

@@ V Experimental Results and Discussion
Our novel framework ChIRAAG, was evaluated using the designs provided in the open-source project OpenTitan [10]. We have performed our experiments on seven designs of the OpenTitan project, which includes RV Timer, PattGen, GPIO, HMAC, ROM Ctrl, sram_ctrl and adc_ctrl.
@@
ChIRAAG generated the correct assertion using LLM within a few automatic prompting processes for each design. Table II shows the performance report for LLM-based assertion generation. The first column, “Module” shows different designs we have taken. The third column shows the number of SV assertions generated by LLM for each design, and the second column, “OT Assert.”, shows the number of assertions provided by the RTL in OpenTitan. “Iterations to Conversion” denotes the number of iterations required to re-prompt the LLM to generate the correct assertions that satisfy all the test cases. “VCS Sim. Time” shows the time taken by the simulation process. While the “ChIRAAG Time” is the SVA generation time taken by our framework.
ChIRAAG successfully generated all of the OpenTitan assertions for each design. Interestingly, ChIRAAG generates more assertions than that was provided in OpenTitan. For example, it generates 12 assertions for HMAC while only four assertions are provided in OpenTitan. Importantly, ChIRAAG-generated assertions are important and cover some aspects of the design that were missed by the OpenTitan assertions. For example, Assertion 3:
@@
assert @(posedge clk) disable iff (rst_n)
(msg_fifo_req && hmac_en) |-> ##1 reg_hash_start)
else \$fatal("Assertion failed");
is crucial to check the HMAC process start property. The HMAC process should start whenever there is a request in the message FIFO and the HMAC signal is enabled. That is, if ‘msg_fifo_req’ and ‘hmac_en’ are true, then the signal ‘req_hash_start’ must be true on the next rising edge of the clock unless a reset (‘rst_n’) is active. If this condition is violated, a fatal error message is triggered. While in the ROM Controller, the additional important SVA that ChIRAAG generated besides the ones that are provided in OpenTitan assertions are:
Assertion 4:
assert @(posedge clk) (checker_done && (q !== Checking && q !== Done’)) |-> ##1 (q $=$ Invalid’);
Assertion 5:
assert @(posedge clk) (counter_done |-> (q !== Reading Low && q_next !== Invalid’)
Assertion 6:
assert @(posedge clk) (((q !== Invalid’) && rst_ni && kmac_rom_last_o) |-> ##1 kmac_rom_vld_o)
These assertions verify some of the essential properties: Assertion 4 checks if ‘checker_done’ is true when the system state (‘q’) is not in the ‘Checking’ or ‘Done’ states. If the condition is violated, the assertion requires that in the next clock cycle, the state ‘q’ must be set to invalid. Assertion 5 states that if the counter_done signal is true, then the state ‘q’ is either not ‘Reading Low’ in the current state or does not become ‘Invalid’ in the next state. While, Assertion 6 states that if the current state ‘q’ is not ‘Invalid’ and both the reset condition ‘rst_ni’ and the signal ‘kmac_rom_last_o’ are true, then in the next clock cycle, the signal ‘kmac_rom_vld_o’ should be true.
TABLE II: Performance Report
Module OT Assert. LLM Assert. Iteration to Conversion VCS Sim. Time ChIRAAG Time
RV Timer 0 9 9 80ns 7.34s
PattGen 0 7 5 120ns 8.35s
GPIO 0 6 10 105ns 9.24s
HMAC 4 12 15 180ns 8.29s
ROM Ctrl 6 14 16 110ns 9.42s
sram_ctrl 0 11 3 160ns 8.34s
adc_ctrl 5 7 5 95ns 7.29s
Also, generating assertions traditionally (manually) is a hectic and time-consuming process. Manually for each design, it would take many hours. Our proposed framework ChIRAAG, does the same work in less than 15 seconds for each design. This shows the effectiveness of ChIRAAG in a faster assertion generation process than the traditional method.
While ChIRAAG can rectify syntactic and simulation errors, it can also detect potential bugs in the implementation. To demonstrate such a scenario, we have considered a buggy implementation of a 4-bit full-adder, where xor is replaced by an or operation. Initially, formatted specifications were fed to ChIRAAG to generate the SVA. ChIRAAG correctly generated two assertions listed as Assertion 7 and Assertion 8, for checking the carry-out generated and the other for checking the sum bits generated.
Assertion 7:
@@
Even though the above assertions are correct, the testcase fails due to the wrong output. This testcase failure log is shown in Fig. 4.
Figure 4: Failure log of full-adder
After correcting the implementation to xor from or, the SVA could pass all the tests, as shown in Fig. 5. This shows that our automated framework can also detect bugs in design implementation.
Figure 5: Success log of full-adder
This shows the usefulness of ChIRAAG in generating SVA from natural language specifications. The basic details of the formatted specification, along with the codebase, simulation files, logs, and prompts, are accessible on our GitHub repository  [^1] Github link will be available upon acceptance.
@@ VI Conclusion
In this work, we have developed ChIRAAG that can generate SV assertion from natural language specification using an LLM as an assertion generator. Primarily, unstructured design specifications are transformed into a structured manner and given to the LLM as an input prompt to generate an SVA suite. The prompting technique is crucial in generating correct assertions. The performance of ChIRAAG on OpenTitan design shows that an LLM is a good starting point to assist an engineer in assertion generation. We noticed only 33% of the generated SVA requires refinement upon iterative prompting. From our observations, it is clear that LLM can be used to produce SV assertion, provided that the LLM will be fed with structured, unambiguous information. We performed our experiment with a gd-LLM; however, we believe a domain-specific LLM for FPV should provide even more accurate assertions in fewer iterations. Although the assertions are correct in satisfying the intended behavior captured by the testbench, it is not possible to confirm 100% functional coverage of design intent. Our future goal will be to check the capability of LLM to address the “consistency” and “completeness” issues in ABV.

### v3

@@ V Experimental Results and Discussion
Our novel framework, ChIRAAG, was evaluated using designs provided in the OpenTitan [5] project. We performed our experiments on six designs as shown in Table II.
@@
ChIRAAG generated the correct assertion using LLM within a few automatic prompting processes for each design. Table II shows the performance report for LLM-based assertion generation. The first column, “Module” shows different designs we have taken. The third column, “LLM Assert.”, shows the number of SV assertions generated by LLM for each design, and the second column, “OT Assert.”, shows the number of assertions provided in the RTL of OpenTitan. “#Prompts” denotes the number of prompts required to generate the correct assertions that satisfy all the test cases. “VCS Sim. Time” shows the time taken by the simulation process. While the “SVA Generation Time” is the initial assertion generation time taken by our framework.
ChIRAAG successfully generated all of the OpenTitan assertions for each design. Interestingly, ChIRAAG generates more assertions than that was provided in OpenTitan. Importantly, ChIRAAG-generated assertions are important and satisfy some aspects of the design that were missed by the OpenTitan assertions. Our proposed framework, ChIRAAG, generates the assertions for each design in less than 15 seconds. This shows the effectiveness of ChIRAAG in a faster assertion generation process than the traditional method.
TABLE II: Performance Report of ChIRAAG
Module OT Assert. LLM Assert. #Prompts VCS Sim. Time SVA Generation Time
RV Timer 0 11 12 80ns 6.34s
PattGen 0 9 9 110ns 9.45s
GPIO 0 6 8 190ns 12.34s
ROM Ctrl 6 11 14 250ns 14.34s
sram_ctrl 0 14 8 100ns 10.23s
adc_ctrl 5 8 9 460ns 7.56s
While ChIRAAG can rectify syntactic and simulation errors using the log of a simulator like VCS, it can also detect potential bugs in the implementation. To demonstrate such a scenario, we have considered a buggy implementation of a 4-bit full-adder, where xor is replaced by an or operation. Initially, formatted specifications were fed to ChIRAAG to generate the SVA. ChIRAAG correctly generated two assertions listed as Assertion 3 and Assertion 4, for checking the carry-out generated and the other for checking the sum bits generated.
@@
Even though the above assertions are correct, the testcase fails due to the wrong output. After correcting the implementation to xor from or, the SVA could pass all the tests, This shows that our automated framework can also detect bugs in design implementation. This shows the usefulness of ChIRAAG in generating SVA from natural language specifications. The basic details of the formatted specification, along with the codebase, simulation files, logs, and prompts, are accessible on our GitHub repository  [^1] https://github.com/karthikmaddala/ChIRAAG.
@@ VI Conclusion
In this work, we have developed ChIRAAG that can generate SVA from the formatted specification using LLM. Primarily, unstructured design specifications are transformed into a structured manner and given to the LLM as an input prompt to generate an SVA suite. The prompting is crucial in generating correct assertions. The RTL implementation is only used to simulate and identify syntax and simulation errors in generated assertions. The performance of ChIRAAG on OpenTitan design shows that an LLM is a good starting point to assist an engineer in assertion generation. We noticed only 27% of the generated SVA require refinement upon iterative prompting. We performed our experiment with a general-domain LLM; however, we believe a domain-specific LLM for ABV should provide even more accurate assertions in fewer iterations. In our work, we ensure that the generated assertions are syntactically and semantically correct. However, how meaningful and complete these are, in the context of a design needs to be checked by experts. It is not possible to confirm 100% functional coverage of design intent by the generated assertions. Our future goal will be to check the capability of LLM to address the “consistency” and “completeness” issues in ABV.